OpenStudy (alexatiger):

Hey guys, I need serious help with this one. For the law of cosines (trying to find side "b") the problem looks like: b^2= 5^2+ 3^2 -2*5*3 cos 90 degrees. I keep getting an odd number, I need help solving this problem and finding b.

OpenStudy (phi):

if you have a 90 degree angle, you have a right triangle. so the problem should reduce to pythagoras. It almost looks like a 3,4,5 right triangle, but the 5 would be the longest side.

OpenStudy (alexatiger):

OH! See, this is my problem and example, I'm supposed to make all of it. Would it help if I changed the 5 to 4 so the hypotenuse can be 5?

OpenStudy (phi):

it would make the numbers nicer. as you (should!) know, cos 90º is 0 so your equation up above simplifies to b^2 = 5^2 + 3^2 b = sqrt(34) if you use 4 instead of 5, you would get b^2 = 4^2+9^2 = 25 b= sqr(25) = 5

OpenStudy (alexatiger):

So the -2ac*cos(90) just disappears?

OpenStudy (phi):

yes, because cos 90 = 0

OpenStudy (alexatiger):

Thank you SO MUCH!!